# Enhancing Hemodynamic Parameter Estimations: Nonlinear Blood Behavior in 4D Flow MRI

### v2

@@ 2.2 On the choice for the shear stress blood behavior
In this work, we will estimate hemodynamic parameters using a Newtonian and a purely viscous shear-thinning blood behavior summarized below:
@@
    Newtonian fluid behavior: This is the gold standard in hemodynamics simulations and assumes that the shear stresses are proportional to the rate of change of the fluid’s velocity vector. While this assumption is widely accepted for large blood vessels where the Reynolds number is also large, its applicability for systolic and diastolic conditions has not yet been thoroughly assessed using modern 4D-flow MRI. The model for a Newtonian fluid involves solving equations (1), where the shear stress is defined as $\tau=2\mu_{\text{N}}\dot{\boldsymbol{\varepsilon}}$, with $\mu_{\text{N}}$ representing the constant viscosity of blood, and $\dot{\boldsymbol{\varepsilon}}$ denoting the symmetric velocity gradient, which satisfies $2\dot{\boldsymbol{\varepsilon}}=\nabla\boldsymbol{u}+\nabla^{T}\boldsymbol{u}$.
@@
    Power-law fluid behavior: Blood is a purely viscous shear-thinning fluid. It can be modeled using a generalized Newtonian law that accounts for strain-rate-dependent viscosity. The power-law model is a two parameters constitutive law that relates the shear stresses with the rate of change of the fluid’s velocity vector as:
    $\tau=\mu_{\text{PL}}\dot{\boldsymbol{\varepsilon}},\quad\text{with }~{}\mu_{\text{PL}}=m\dot{\gamma}^{n-1}$ (3)
@@
2.3 Signal model for 4D Flow images
The generation of 4D Flow MR images was based on the blood velocity $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ obtained from CFD simulations. The process involved evaluating the signal equation, which depends on the magnetization at thermal equilibrium $M_{0}(\boldsymbol{r})$, velocity encoding sensitivity (VENC), blood velocity $u_{a}$ (with $a=x,y,z$), and the relaxation time of the blood $T_{2}^{*}(\boldsymbol{r})$. The signal equation is expressed as follows:
@@ 2.3 Signal model for 4D Flow images
$M(\boldsymbol{\boldsymbol{r}},t)$ $=M_{0}(\boldsymbol{r})e^{-\boldsymbol{i}\pi/\text{VENC}u_{a}(\boldsymbol{r})}e^{-t/T_{2}^{*}(\boldsymbol{r})},$ (4b)
In this formulation, $p(\boldsymbol{r})$ represents the slice profile, and $\boldsymbol{k}(t)$ denotes the Fourier space trajectory during signal acquisition, which varies based on the chosen imaging sequence. The integration domain $B$ in the integral refers to the object being measured, in our case, the aorta. Careful consideration of scanner hardware limitations is required when determining the time $t$ at which each $\boldsymbol{k}(t)$ location is measured. Factors such as the velocity, phase-encoding, and readout gradients, which depend on the maximum slew rate and amplitude of the gradient system, influence the echo time. Additionally, the readout gradient is affected by the sampling frequency of the ADC, which impacts not only the timing of each $k$-space measurement but also the occurrence of certain artifacts.
During signal measurement, the positions of spins change due to blood dynamics. This motion is accounted for using a first-order approximation of the spins’ positions $\boldsymbol{r}(t)$ around $t_{k}$:
$\boldsymbol{r}(t)=\boldsymbol{r}(t_{k})+\boldsymbol{u}(t_{k})(t-t_{k})+\mathcal{O}(t^{2}),$ (5)
@@
No field inhomogeneities were considered during the evaluation of the signal equation, and the integral was calculated for each $k$-space location using the Finite Element Method (FEM). All the steps described here were implemented in a branch of the PyMRStrain library (github.com/hmella/pymrstrain/tree/4dflow), which has previously been used to generate synthetic strain phantoms [38, 39, 40]. The simulation workflow described in this section is illustrated in Figure 1.
@@
2.4 Estimation of power-law parameters for custom Hct values and fitted-Newtonian viscosities
Viscosity measurements at 37o for different Hct values of 16%, 33%, 43%, 57%, and 70% across a shear rate range of 12 to 123 s-1 were obtained from [18]. Since the experimental data were not measured at the same shear rates, the parameters $m$ and $n$ in the power-law model described in (3) were fitted for the values provided in the article within a shear rate range of 10 to 130 s-1. One-dimensional interpolations between these fits were then performed to generate synthetic experimental measurements at shear rates of 12.3, 24.4, 61.8, and 123.7 s-1 (mean shear rate across experimentally measured Hct values). These interpolated values were subsequently used to fit the power-law model over the same shear rate range described previously.
@@ 2.4 Estimation of power-law parameters for custom Hct values and fitted-Newtonian viscosities
Figure 2: Viscosities obtained from Wells et al. [18]. (a) Experimental measurements and the adjusted power-law viscosities ($\mu_{\mathrm{PL}}$) derived from the data. (b) Synthetic measurements (obtained using one-dimensional linear interpolation), along with Newtonian fits ($\mu_{\mathrm{NF}}$) and power-law fits ($\mu_{\mathrm{PL}}$) for intermediate Hct levels not reported in [18]. Colors represents different Hct levels.
The power-law fit was performed using weighted least squares to account for unequal variances across shear rates and to avoid unwanted fitting biases. The Newtonian fit was carried out by estimating the mean value of the synthetic measurements, effectively solving a least squares problem for a constant model. The resulting viscosities are shown in Figure 2.
@@
With this in mind, it is important to note that three different viscosities were considered in this investigation:
  • 1.
@@
  • 3.
    $\mu_{\text{PL}}$: The non-linear viscosity estimated using the power-law model given in (3) and fitted to Hct values.
All of these viscosities are listed in Table 1 for Hct values of 20, 32.5, 45, 57.5, and 70, which were used in all experiments involving CFD simulations and synthetic images.
@@ 3 Experiments
Table 1: Viscosities used for CFD simulations and the estimation of hemodynamic parameters from synthetic and in-vivo 4D Flow images were obtained from [18] using the procedure described in Section 2.4. The parameters $m$ and $n$ were used to calculate $\mu_{\text{PL}}$ by evaluating Equation (3).
Hematocrit (%) Newtonian Power-law11footnotemark: 1 Newtonian fitted
@@
28.2 3.0, 3.5, 4.0, 4.5 1.3604, 0.6506 -
35.2 1.8313, 0.6655 -
40.2 2.0307, 0.7143 -
46.6 2.6352, 0.7015 -
50.1 3.1224, 0.6759 -
3.1 Computational fluid dynamics for flow simulations
For all CFD experiments, the computational domain was set as a tetrahedral mesh of a patient-specific aorta model (see Figure 1), obtained from the Vascular Model Repository [2, 41, 42, 43]. The mesh consists of 127,131 vertices and 719,419 tetrahedrons, and it has been refined near the wall to account for boundary layer effects and to accurately represent spatial gradients (see Figure 3).
The governing equations of the flow given in (1) were solved using FEM with piece-wise linear Lagrange elements for $\boldsymbol{u}$ and $p$. To overcome the well-known inf-sup constraints [44], a Brezzi-Pitkäranta stabilization term was added to the formulation [45], alongside a streamline upwind term to ensure stability for highly convective flows [46, 47]. Similar stabilization techniques in non-Newtonian fluids have been tested and analyzed in detail in [48, 49]. Additionally, the back-flow stabilization method proposed in [50, 51] was incorporated into all the simulations.
The inlet boundary condition, as defined in (1c), is represented by $g(x,t)=u_{\text{max}}(t)\mathcal{S}(x)$. Here, $u_{\text{max}}$ is obtained from phase-contrast MRI measurements [2]. The function $\mathcal{S}$ characterizes the spatial distribution of the Dirichlet condition, resembling a laminar parabolic fluid profile. Both $u_{\text{max}}$ and $\mathcal{S}$ are shown in Figure 3.
@@ 3.1 Computational fluid dynamics for flow simulations
We employed a second-order finite difference formula for temporal discretization with a time step size of $\Delta t=9.37\times 10^{-4}$ s (equivalent to 1000 steps per cardiac cycle). The simulation covered two cardiac cycles, with only the last cycle considered in the analysis to avoid nonphysical solutions.
The non-linear terms in (1), i.e., the convection and the fluid constitutive, were linearized using a Picard fixed point approach, with an $H^{1}\times L^{2}$ error tolerance of $\delta_{p}=10^{-4}$. The coupling with the Windkessel model is done semi-implicitly, using a 4th-order Runge-Kutta scheme for the ODE.
We use the software MAD [52] to assemble the discrete equations and invert the linear systems. MAD is built upon the linear algebra library PETSc [53], and the computations are done with up to 72 CPU cores in parallel. A summary of the numerical simulation parameters is given in Table 2.
@@ 3.2 Generation of synthetic 4D Flow images
To emulate realistic imaging conditions, Gaussian noise was introduced with a standard deviation set at 5.2% of the maximum amplitude of the complex signal’s magnitude, achieving an SNR of 13.5, which aligns with the average SNR reported for the ascending and descending aorta at 1.5T [57]. Noise was added to both the real and imaginary channels.
Finally, the MRI signals generated from CFD simulations, using the power-law model and constant viscosities adjusted for Hct, were reconstructed into 4D Flow images by performing the inverse Fourier transform of the signal, as defined in equation (4).
@@ 3.3 Computation and Comparison of Hemodynamic Parameters
In all the experiments conducted in this work, the primary focus is on comparing WSS, OSI, and $\dot{E}_{L}$ [8, 25] parameters due to their dependence on blood viscosity. These parameters were estimated as follows:
$\text{WSS}(\boldsymbol{x},t)$ $=\lVert\boldsymbol{l}\times\left\{\left(2\mu\dot{\boldsymbol{\varepsilon}}\boldsymbol{l}\right)\times\boldsymbol{l}\right\}\rVert_{2},$ reported in Pascals (6a)
@@
$\dot{E}_{L}(\boldsymbol{x},t)$ $=2\mu\left(\dot{\boldsymbol{\varepsilon}}-\frac{2}{3}\nabla\cdot\boldsymbol{u}\textbf{I}\right):\left(\dot{\boldsymbol{\varepsilon}}-\frac{2}{3}\nabla\cdot\boldsymbol{u}\textbf{I}\right)V(\boldsymbol{x}),$ reported in micro-Watts (6c)
Here, $\boldsymbol{l}$ denotes a unit inward surface normal vector, $2\mu\dot{\boldsymbol{\varepsilon}}\boldsymbol{l}$ represents the shear stress, $T$ is the duration of the cardiac cycle, $V(\boldsymbol{x})$ is the Voronoi nodal volume distribution on the finite element mesh, and I is the identity matrix. Additionally, $\lVert\cdot\rVert_{2}$ denotes the standard Euclidean vector norm. In the non-Newtonian case, the apparent viscosity varies both spatially and temporally, and this variation is accounted for by using $\mu_{\text{PL}}$ in the relevant expressions.
These three biomarkers were chosen because of their direct dependence on viscosity. The boundary layer near the vessel wall, influenced by blood rheology, can undergo thinning or thickening, which affects velocity gradients and, consequently, viscosity. In this context, evaluating WSS is crucial. OSI measures temporal changes in shear stress caused by variations in viscosity and flow dynamics, aspects not captured by $\boldsymbol{t}_{\text{WSS}}$. Finally, $\dot{E}_{L}$ quantifies how energy dissipates within the flow due to viscous fluid interactions, which are also sensitive to changes in viscosity.
To compare these three parameters, the aorta was divided into four segments: the ascending aorta (AAo, between the Valsalva level and the brachiocephalic trunk), the aortic arch (AArch, between the brachiocephalic trunk and the isthmus level), the proximal descending aorta (pDAo, between the isthmus level and the Valsalva level), and the distal descending aorta (dDAo, between the Valsalva level and the diaphragmatic level). This division into segments is not arbitrary and follows the anatomical references used in previous studies [5, 59, 60, 61, 1]. A schematic of these segments is shown in Figure 4. The mean values of WSS, OSI, and $\dot{E}_{L}$ were then estimated for each segment to facilitate comparisons. In line with current practices [3, 62, 60, 63, 64], the supra-aortic branches (brachiocephalic, carotid, and subclavian arteries) were not included in the analysis due to the small number of voxels relative to artery diameters in this region, which affects the quantification of hemodynamic parameters.
The three quantities specified in (6) were estimated using the 4D Flow Matlab Toolbox [65], which employs a previously validated finite element scheme [66] to compute various hemodynamic and geometrical parameters from 4D Flow MRI data [1].

### v3

@@ 2.2 On the choice for the shear stress blood behavior
In this work, blood rheology was modeled using both a Newtonian and a non-Newtonian purely viscous shear-thinning blood behavior. A Power-law curve was chosen to characterize the blood rheology, as it is both simple (i.e., depends only on two parameters) and flexible enough to reproduce experimental data at physiological ranges. In addition, there are accepted correlations for power-law curves and the Hct level, which is directly measurable through a blood test (a necessary condition for in-vivo experiments and always available in the context of a cardiovascular MRI exam). Consequently, no additional efforts are required to define the patient’s blood rheology. In the following paragraphs, both models are summarized:
@@
    Newtonian blood behavior: This is the gold standard in hemodynamics simulations and assumes that the shear stresses are proportional to the rate of change of the fluid’s velocity vector. While this assumption is widely accepted for large blood vessels where the Reynolds number is also large, its applicability for systolic and diastolic conditions has not yet been thoroughly assessed using modern 4D Flow MRI. The model for a Newtonian fluid involves solving equations (2.1), where the shear stress is defined as $\tau=2\mu_{\text{N}}\dot{\boldsymbol{\varepsilon}}$, with $\mu_{\text{N}}$ representing the constant viscosity of blood, and $\dot{\boldsymbol{\varepsilon}}$ denoting the symmetric velocity gradient, which satisfies $2\dot{\boldsymbol{\varepsilon}}=\nabla\boldsymbol{u}+\nabla^{T}\boldsymbol{u}$.
@@
    Power-law blood behavior: Blood is a fluid that can be characterized by shear-thinning behavior and a plastic component associated with its yield stress [23]. The power-law model is a two-parameter constitutive law that relates shear stress to the rate of change of the fluid’s velocity vector, without accounting for the plastic component of blood, as:
    $\tau=\mu_{\text{PL}}\dot{\boldsymbol{\varepsilon}},\quad\text{with }~{}\mu_{\text{PL}}=m\dot{\gamma}^{n-1}$ (2.3)
@@
It is important to note that inhomogeneous blood phenomena were not considered in the selection of the rheological model. For instance, the Fahraeus effect [41], plasma skimming [54], and the spatial distribution of red blood cells [55] (among others) were excluded either because they do not occur in vessels with large diameters, such as the aorta and its branches, or because the rheology model used does not account for spatially varying Hct (as this is beyond the scope of this investigation). Additionally, in MRI studies (which are describe later), it is uncommon to evaluate hemodynamic parameters in aortic branches due to resolution limitations.
2.3 Signal model for 4D Flow images and generation of synthetic MR images
The generation of 4D Flow MR images used as input the blood velocity $\boldsymbol{u}=(u_{x},u_{y},u_{z})$ obtained from CFD simulations. The image generation process involved evaluating the signal equation, which depends on the magnetization at thermal equilibrium $M_{0}(\boldsymbol{r})$, velocity encoding sensitivity (VENC), blood velocity $u_{a}$ (with $a=x,y,z$), and the relaxation time of the blood $T_{2}^{*}(\boldsymbol{r})$. The signal equation is expressed as follows:
@@ 2.3 Signal model for 4D Flow images and generation of synthetic MR images
$M(\boldsymbol{\boldsymbol{r}},t)$ $=M_{0}(\boldsymbol{r})e^{-\boldsymbol{i}\pi/\text{VENC}u_{a}(\boldsymbol{r})}e^{-t/T_{2}^{*}(\boldsymbol{r})},$ (2.4b)
In this formulation, $p(\boldsymbol{r})$ represents the slice profile, and $\boldsymbol{k}(t)$ denotes the Fourier space trajectory during signal acquisition, which varies based on the chosen imaging sequence. The integration domain $B$ in the integral refers to the object being measured, in our case, the aorta. Careful consideration of hardware limitations of the scanner are required when determining the time $t$ at which each $\boldsymbol{k}(t)$ location is measured. Factors such as the velocity-encoding gradients, phase-encoding gradients, and readout gradients, which depend on the maximum slew rate and amplitude of the gradient system, influence the echo time. Additionally, the readout gradients are affected by the sampling frequency of the ADC, which impacts not only the timing of each $k$-space measurement but also the occurrence of certain artifacts.
During signal measurement, the positions of spins change due to blood dynamics. This motion was accounted for in the generation of synthetic images using a first-order approximation of the spins’ positions $\boldsymbol{r}(t)$ around $t_{k}$:
$\boldsymbol{r}(t)=\boldsymbol{r}(t_{k})+\boldsymbol{u}(t_{k})(t-t_{k})+\mathcal{O}(t^{2}),$ (2.5)
@@
No field inhomogeneities were considered during the evaluation of the signal equation, and the integral was calculated for each $k$-space location using the Finite Elements Method (FEM). All the steps described here were implemented in the Finite Element MRI (FEelMRI) library (github.com/hmella/FEelMRI), which has previously been used to generate synthetic strain phantoms [56, 57, 58], and was used to generate the synthetic images in this study. The simulation workflow described in this section is illustrated in Figure 1.
@@
2.4 Estimation of power-law parameters for custom Hct values and Newtonian fitted viscosities
Experimental viscosity measurements at 37o for Hct values of 16%, 33%, 43%, 57%, and 70%, across a shear rate range of 12 to 123 s-1, were obtained from [24] (crosses in Figure 2a). The parameters $m$ and $n$ in the power-law model described in (2.3) were estimated from these experimental data using a weighted least-squares fitting method to avoid data over-representation.To determine $m$ and $n$ values for any Hct within the range 16% to 70%, we assumed that the fitted power-law viscosity curves (see Figure 2a) varied linearly between the given Hct values.
The procedure for obtaining intermediate Hct viscosity curves was as follows:
  1. 1.
    For a given intermediate Hct value, we identified the two fitted curves between which this value lies.
  2. 2.
    Intermediate synthetic experimental measurements (crosses in Figure 2b), at the same shear rates as the original experimental data, were generated using linear interpolation between the two curves.
  3. 3.
    The intermediate synthetic experimental measurements were then used to fit the parameters $m$ and $n$ using a weighted least-squares fitting method (curves in Figure 2b).
It is important to note that our decision to follow this procedure was based on the fact that we could not assume $n$ varies linearly across Hct, as no clear trend was observed between these parameters (see Table 1). This procedure was applied to obtain the power-law parameters at Hct values of 20, 32.5, 45, 57.5, and 70% (see Table 1).
For the Newtonian viscosities fitted to Hct values, an analytical and physically consistent method designed to preserve the average shear stress within a specific shear rate range for a given Hct value was adopted. This method can be directly applied to a power-law viscosity curve to compute a corresponding Newtonian viscosity, as follows:
[…]
where $\mu_{\text{PL}}$ represents the power-law viscosity obtained using (2.3), and $\mu_{\text{NF}}$ denotes the Newtonian fitted viscosity. However, this approach yields elevated Newtonian viscosities when constrained to the limited shear rate range (12 to 123 s-1) of the experimental measurements (see Newtonian fit 1 in Table 1). To mitigate this issue, a second approach extends the shear rate range of the synthetic experimental measurements to include values where blood exhibits Newtonian behavior. For instance, expanding the range to 0–2800 s-1 results in lower Newtonian viscosities (see Table 1), which, for a healthy Hct of 45%, align with values commonly reported in the literature (see Newtonian fit 2 in Table 1).
A third approach involves adopting the model used in [46], which is derived by fitting a 5th-order polynomial to viscosity measurements from [59]. This method produces a Newtonian viscosity model that depends on Hct and can be applied to any Hct value. Additionally, it yields viscosities within the range reported in the literature (see Newtonian fit 3 in Table 1). However, the conditions under which these measurements were obtained remain unclear.
We chose the second approach (hereafter referred to as Newtonian fitted viscosities) for all CFD simulations of Hct-varying shear-thinning blood because, to the best of our judgment, it represents the most physically consistent translation between both models given the available data. This method ensures alignment with established values in the literature for linear models (at least for healthy Hct) while also accounting for a wide range of shear rate regimes, such as those found in the aorta [60, 61, 44], including highly convective and even pathological scenarios.
@@ 2.4 Estimation of power-law parameters for custom Hct values and Newtonian fitted viscosities
Figure 2: Viscosities obtained from Wells et al. [24]. (a) Experimental measurements and the adjusted power-law viscosities ($\mu_{\mathrm{PL}}$) derived from the data. (b) Synthetic measurements (obtained using one-dimensional linear interpolation) and power-law fits ($\mu_{\mathrm{PL}}$) for intermediate Hct levels not reported in [24]. Colors represents different Hct levels and, for each level, the coefficient of determination ($R^{2}$) and root-mean-squared error (RMSE) are reported.
@@
Based on the above, it is important to note that three different viscosities were considered in this investigation for the generation of synthetic 4D Flow MR images:
@@
  • $\mu_{\text{PL}}$: The non-linear viscosity estimated using the power-law model given in (2.3) and fitted to Hct values.
All of these viscosities are listed in Table 1 for Hct values of 20, 32.5, 45, 57.5, and 70, which were used in all experiments involving CFD simulations for the generation of synthetic 4D Flow MR images and the estimation of hemodynamic parameters from these data.
@@ 3 Experiments
Table 1: Viscosities used for CFD simulations and the estimation of hemodynamic parameters from synthetic and in-vivo 4D Flow images were obtained from [24] using the procedure described in Section 2.4. The parameters $m$ and $n$ are used to calculate $\mu_{\text{PL}}$ by evaluating Equation (2.3). Emphasized column denotes the Newtonian viscosity fit used for the CFD simulations in Experiment 1.
Hematocrit (%) Newtonian Power-law11footnotemark: 1 Newtonian fit 1 Newtonian fit 2 Newtonian fit 3
@@
28.2 3.0, 3.5, 4.0, 4.5 1.36, 0.65 - - -
35.2 1.83, 0.67 - - -
40.2 2.03, 0.71 - - -
46.6 2.64, 0.70 - - -
50.1 3.12, 0.68 - - -
3.1 Computational fluid dynamics for blood flow simulations
For all CFD experiments, the computational domain was set as a tetrahedral mesh of a patient-specific aorta model (see Figure 1), obtained from the Vascular Model Repository [8, 62, 63, 64]. The mesh comprises of 127,131 vertices and 719,419 tetrahedrons and was refined near the wall to account for boundary layer effects and to accurately represent spatial gradients (see Figure 3).
The governing equations of the flow given in (2.1) were solved using FEM with piece-wise linear Lagrange elements for $\boldsymbol{u}$ and $p$. To overcome the well-known inf-sup constraints [65], a Brezzi-Pitkäranta stabilization term was added to the formulation [66], alongside a streamline upwind term to ensure stability for highly convective flows [67, 68]. Similar stabilization techniques in non-Newtonian fluids have been tested and analyzed in detail in [69, 70]. Additionally, the back-flow stabilization method proposed in [71, 72] was incorporated into all the simulations.
The inlet boundary condition, as defined in Equation (2.1c), is represented by $g(x,t)=u_{\text{max}}(t)\mathcal{S}(x)$. Here, $u_{\text{max}}$ was obtained from phase-contrast MRI measurements [8] and the function $\mathcal{S}$ characterizes the spatial distribution of the Dirichlet condition, resembling a laminar parabolic fluid profile. Both $u_{\text{max}}$ and $\mathcal{S}$ are shown in Figure 3.
@@ 3.1 Computational fluid dynamics for blood flow simulations
A second-order finite difference formula with a time step size of $\Delta t=9.37\times 10^{-4}$ s (equivalent to 1000 steps per cardiac cycle) was used for temporal discretization. Two cardiac cycles were simulated for each experiment, and only the last cycle was considered in the analysis to avoid nonphysical solutions.
The non-linear terms in (2.1), i.e., the convection and the fluid constitutive, were linearized using a Picard fixed point approach, with an $H^{1}\times L^{2}$ error tolerance of $\delta_{p}=10^{-4}$. The coupling with the Windkessel model was done semi-implicitly, using a 4th-order Runge-Kutta scheme for the ODE.
The software MAD [73] was used to assemble the discrete equations and invert the linear systems. MAD is built upon the linear algebra library PETSc [74], and the computations are done with up to 72 CPU cores in parallel. A summary of the numerical simulation parameters is given in Table 2.
@@ 3.2 Generation of synthetic 4D Flow images
To emulate realistic imaging conditions, independent and uncorrelated Gaussian noise was added to both the real and imaginary components of the signal. The noise standard deviation was set at 5.2% of the maximum amplitude of the complex signal’s magnitude, resulting in a signal-to-noise ratio (SNR) of 13.5. This value is consistent with the average SNR reported for the ascending and descending aorta at 1.5T [77].
Finally, the MRI signals generated from CFD simulations, using the power-law model and constant viscosities adjusted for Hct, were reconstructed into 4D Flow images by performing the inverse Fourier transform of the signal, as defined in equation (2.4).
@@ 3.3 Computation and Comparison of Hemodynamic Parameters
In all the experiments conducted in this work, the primary focus was on comparing WSS, OSI, and $\dot{E}_{L}$ [14, 46] parameters due to their dependence on blood viscosity. These parameters were estimated as follows:
$\text{WSS}(\boldsymbol{x},t)$ $=\lVert\boldsymbol{l}\times\left\{\left(2\mu\dot{\boldsymbol{\varepsilon}}\boldsymbol{l}\right)\times\boldsymbol{l}\right\}\rVert_{2},$ reported in Pascals (3.1a)
@@
$\dot{E}_{L}(\boldsymbol{x},t)$ $=2\mu\left(\dot{\boldsymbol{\varepsilon}}-\frac{2}{3}\nabla\cdot\boldsymbol{u}\textbf{I}\right):\left(\dot{\boldsymbol{\varepsilon}}-\frac{2}{3}\nabla\cdot\boldsymbol{u}\textbf{I}\right)V(\boldsymbol{x}),$ reported in micro-Watts (3.1c)
Here, $\boldsymbol{l}$ denotes a unit inward normal vector to the surface, $2\mu\dot{\boldsymbol{\varepsilon}}\boldsymbol{l}$ represents the shear stress, $T$ is the duration of the cardiac cycle, $V(\boldsymbol{x})$ is the Voronoi nodal volume distribution on the finite element mesh, and I is the identity matrix. Additionally, $\lVert\cdot\rVert_{2}$ denotes the standard Euclidean vector norm. In the non-Newtonian case, the apparent viscosity varies both spatially and temporally, and this variation was accounted for by using $\mu_{\text{PL}}$ in the relevant expressions.
These three biomarkers were chosen because of their direct dependence on viscosity. Although near-wall velocity gradients are elevated (suggesting that blood behaves more like a Newtonian fluid in these regions) the pronounced shear-thinning behavior away from the aortic walls (where shear rates are lower), combined with the occurrence of vortex shedding, affects the entire velocity profile and consequently influences the velocity gradients at the wall. In this context, evaluating WSS is crucial. OSI measures temporal changes in shear stress caused by variations in viscosity and flow dynamics, aspects not captured by $\boldsymbol{t}_{\text{WSS}}$. Finally, $\dot{E}_{L}$ is a volumetric parameter quantifying how energy dissipates within the flow due to viscous fluid interactions, which are also sensitive to changes in viscosity, allowing for evaluation away from the vessel walls.
@@
To compare WSS, OSI, and $\dot{E}_{L}$, the aorta was divided into four segments: the ascending aorta (AAo, between the Valsalva level and the brachiocephalic trunk), the aortic arch (AArch, between the brachiocephalic trunk and the isthmus level), the proximal descending aorta (pDAo, between the isthmus level and the Valsalva level), and the distal descending aorta (dDAo, between the Valsalva level and the diaphragmatic level). This segmentation is not arbitrary and follows the anatomical references used in previous studies [11, 79, 80, 81, 6]. A schematic of these segments is shown in Figure 4. The mean values of WSS, OSI, and $\dot{E}_{L}$ are then estimated for each segment to facilitate comparisons. In line with current practices [9, 7, 80, 82, 83], the supra-aortic branches (brachiocephalic, carotid, and subclavian arteries) were not included in the analysis due to the small number of voxels relative to artery diameters in this region, which affects the quantification of hemodynamic parameters.
The three quantities specified in (3.1) were estimated using the 4D Flow Matlab Toolbox [84], which employs a previously validated finite element scheme [85] to compute various hemodynamic and geometrical parameters from 4D Flow MRI data [6]. The complete workflow for estimating hemodynamic parameters from 4D Flow MR images is shown in Figure 5.
Figure 5: Workflow for estimating hemodynamic parameters from velocity measurements obtained using 4D Flow MRI. First, the image segmentation is derived from the magnitude component of the 4D Flow MRI data. Second, a FE mesh is generated by triangulating the segmented surface and constructing tetrahedral elements for the interior volume. Third, blood velocity data obtained from the image phase information is linearly interpolated onto the FE mesh without any smoothing or correction. Fourth, hemodynamic parameters are estimated by computing velocity gradients through a least-squares projection in the FE sense, as described in [85] and implemented in [84]. Finally, mean values and standard deviations of the hemodynamic parameters are calculated for each aortic segment.